OpenStudy (anonymous):

Dividing a polynomial by a monomial: Multivariate? Divide (20z^6u^6-36z^6u^2)/(4z^5u^3)

OpenStudy (lochana):

\[\frac{20z^{6u^6} - 36z^{6u^2}}{4z^{5u^{3}}}\]\[\frac{20z^{6u^6}}{4z^{5u^{3}}} - \frac{36z^{6u^2}}{4z^{5u^{3}}}\]

OpenStudy (lochana):

the key thing here is that you have to subtract exponents when dividing

OpenStudy (anonymous):

I have to simplify it as much as possible

OpenStudy (anonymous):

there are 3 questions that i really cant figure the answer to, i did the rest but the 3 questions are so hard and i need to get them in before 12 tonight.

OpenStudy (lochana):

yes. can you see \[6u^6 \ , 5u^3\] at the first part

OpenStudy (lochana):

mm.. you need to subtract them \[5z^{6u^6 - 5u^3}\]

OpenStudy (lochana):

can you do second one \[6u^2, \ 5u^3\]

OpenStudy (lochana):

well it's \[9z^{6u^2 - 5u^3}\] so finally we get \[5z^{6u^6-5u^3} - 9z^{}6u^2-5u^3\]

OpenStudy (lochana):

\[5z^{6u^6 - 5u^3} - 9z^{6u^2 - 5u^3}\]